Solve the problem.
8). The sum of four times a number and three is the same as the difference of twice the number and eleven. Find the
number.
9). Four times the difference of a number and one is equal to six times the sum of the number and three. Find the
number.

Answer by BurningFlame_marc(77)

Hello!

8.let x=the number
Solution:
4x+3=2x-11
4x=2x-14
2x=-14
x=-7
Final Answer:-7

9.let x=the number
Solution:
4(x-1)=6(x+3)
4x-4=6x+18
4x-6x=18+4
-2x=22
x=-11